# High-dimensional analysis of ridge regression for non-identically distributed data with a variance profile

### v3

@@ 4 Numerical experiments
In this section, we illustrate the results of this paper with numerical experiments. Although we do not have an explicit formula for $T_{p}(z)$, this matrix-valued function satisfies the fixed-point equation (2.3). This allows us to approximate $T_{p}(z)$ numerically using a fixed-point algoritheorem. In this manner, we are thus able to obtain a numerical approximation of the deterministic equivalent $r_{\lambda}^{test}(X_{n})$ of the predictive risk. In this section, we consider several variance profiles that we normalize such that
$\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{p}\frac{\gamma^{2}_{ij}}{p}=1.$
Apart from the constant, quasi doubly stochastic, and piecewise constant variance profiles mentioned earlier, we will use the following examples of variance profiles:
@@
A variance profile referred to as the Berlin Photo for which $\gamma_{ij}$ is equal to the value of the coordinate $(i,j)$ of the pixel of the image depicted in Figure 2. The original image being of size $1250\times 850$, it is rescaled according to the values of $n$ and $p$ used in the numercial experiments thanks to the function resize from the Python module PIL.Image.
@@
The values of the predictive and training risks and their deterministic equivalent are compared for various variance profiles in Figures 3 and 3 with $\lambda$ ranging from 0.1 to 10, $n=400$ and $p=600$. The curves displayed in these figures confirm that $r_{\lambda}^{test}(X_{n})$ and $r_{\lambda}^{train}(X_{n})$ are very accurate estimators of $\hat{r}_{\lambda}^{test}(X_{n})$ and $\hat{r}_{\lambda}^{train}(X_{n})$ in high-dimension since the dashed and solid lines coincide. These variance profiles also provide curves of predictive risks having similar shapes (up to a vertical translation). The curves representing the cases of constant and quasi doubly stochastic variance profile coincide, confirming the comment made in Section 1.1 on the doubly stochastic profile and its similarities with the setting of iid data. Moreover, the minimum of $r_{\lambda}^{test}(X_{n})$ is indeed reached at the optimal value $\lambda_{\ast}$ for any variance profile as shown by Corollary 3.3.
@@
We illustrate the appearance of a double or triple descent phenomenon for several variance profiles in Figure 4 (a), (c) and (e). These figures represent the predictive risk and its approximation by a deterministic equivalent for various values of the ratio $p/n$ with $\lambda=0$. The solid lines depicts $r_{0}^{test}$ whereas the dashed lines represent $\hat{r}_{0}^{test}$. For every variance profile, the solid and dashed lines coincide which confirms that $r_{0}^{test}$ is a relevant deterministic equivalent of $\hat{r}_{0}^{test}$. For the constant and quasi doubly stochastic variance profiles, we observe the well known double descent phenomenon as illustrated by Figure 4 (a) since the curves are increasing for $p/n<1$ and decreasing for $p/n>1$. This is related to Corollary 3.2 which states that the expression of $\lim_{\lambda\rightarrow 0}r_{\lambda}^{test}(X_{n})$ depends upon the value of the ratio $p/n$ with respect to one. Nevertheless, for other variance profiles, the shape of the predictive risk can be very different from one variance profile to another, and it differs from the usual double descent. For some variance profiles a phenomenon of triple descent arises, notably for the piecewise constant and block cases, as shown in Figure 4 (c) and (e). We also remark the appearance of a quadruple descent in the case of the polynomial profile, as shown in Figure 4 (e).
@@
As already remarked at the end of Section 3, it is difficult to analytically explain these phenomenon for an arbitrary variance profile since we do not have an explicit formula for $T_{p}(z)$ in the general case. Nevertheless, as nicely discussed in [SKR+23], the double descent phenomenon is very much related to the distance from zero of the smallest non-negative eigenvalue of $\widehat{\Sigma}_{n}$. Having this eigenvalue close to zero causes double descent. In Figure 4 (b), (d) and (f), we thus represent the smallest non-zero eigenvalue $\tau_{\min}(\widehat{\Sigma}_{n})$ of $\widehat{\Sigma}_{n}$ for many variance profiles as a function of $p/n$. It can be observed that all the curves $p/n\to\tau_{\min}(\widehat{\Sigma}_{n})$ reach their minima at a value close to zero when $p=n$ where the double descent occurs, which is consistent with previous assertions in [SKR+23]. These curves in the constant and quasi doubly stochastic cases behave in a similar same way. However, one can remark from Figure 4 (f) that, for the piecewise constant and polynomial variance profiles, the value of $\tau_{\min}(\widehat{\Sigma}_{n})$ stay much closer to 00 when $p/n$ is between $0.25$ and $3$ than in the constant or quasi doubly stochastic cases. The fact that these curves have a plateau near zero for $p/n\in[0.25,3]$ may be related to the appearance of the triple and quadruple descent phenomenon that is observed in Figure 4 (e).
@@
Figure 5: Study of MNIST dataset whenever $\lambda=0$. (a) Comparison of real and synthetic data with $\alpha=\sigma=1$, $p=784$ and $n$ varying from $78$ to $1568$. The solid lines correspond to the predictive risk while the dashed line corresponds to the deterministic equivalent.(b) Heatmap of the variance of pixels for each class. These heatmaps serve as variance profiles in the mixture model described below.
@@ 5 Conclusion
In this paper, we derive a deterministic equivalent of the DOF and the predictive risk of ridge (less) regression in a high-dimensional framework with a variance profile to handle the setting of non-iid data. We use RMT to determine a deterministic equivalent of the diagonal of the resolvent matrix $Q_{p}(z)$. Our work extends the study of the DOF and the predictive risk from [DW18] and [Bac24] to the case of general variance profile. It appears that also the result of [DW18] and [Bac24] still hold in the case of a quasi doubly stochastic profile. However, there are variance profiles that cause a behavior of the predictive risk of the minimum norm least square estimator that differs from the double descent phenomenon as classically observed in the case of a constant variance profile. Our numerical experiments confirm that our deterministic equivalent accurately estimates the predictive risk in high-dimension. Our results also allow to understand how assuming such a variance profile for the data influences the statistical properties of ridge regression when compared to the standard assumption of iid observations. We hope that our approach on the use of variance profiles may lead to further research works on the statistical analysis of other estimators than ridge regression in more complex models with non-iid data.
@@
In this appendix, we give all the proofs of the results of the paper. We also discuss the link between random matrices with a variance profile and the notion of $\mathcal{R}$-transform in free probability.
@@
[…]
@@ A.2 Derivation of the deterministic equivalent the diagonal of the resolvent
This section is dedicated to prove Theorem 3.1 that aims to provide a deterministic equivalent of $\Delta[Q_{p}(z)]$ that is necessary to obtain asymptotic equivalents of the risks. The proof of this theorem is strongly inspired by results from [HLN07], and it constitutes a key element in the proof of Theorem 3.2. In order to prove Theorem 3.1, we introduce the diagonal matrices $R_{p}(z)=\underset{1\leq j\leq p}{\mathop{\rm diag}\nolimits}(R^{(j)}_{p}(z))$ defined by
$R^{(j)}_{p}(z)=\frac{-1}{z\left((1+(1/n)\operatorname{Tr}[\widetilde{D}_{j}\widetilde{Q}_{n}(z)]\right)}\quad\mbox{for }z\in\mathbb{C}\setminus\mathbb{R}^{+},$
with $\widetilde{Q}_{n}(z)=\Big{(}\widetilde{\Sigma}_{n}-zI_{n}\Big{)}^{-1}$ is the resolvent matrix of $\widetilde{\Sigma}_{n}=\frac{1}{n}X_{n}X_{n}^{\top}$.
@@
Then, Lemma A.1 needs to be combined with the following one in order to prove Theorem 3.1.
@@
Hence, for each compact subset $C\subset\mathbb{C}\setminus\mathbb{R}^{+}$, $f_{p}$ is uniformly bounded on $C$, that is $|f_{p}(z)|\leq\frac{2K}{\delta_{C}}$, where $\delta_{C}$ is the distance between $C$ and $\mathbb{R}^{+}$. Then, by the normal family Theorem [Rud87][Theorem 14.6] there exists a sub-sequence $f_{p_{k}}$ which uniformly converges to $f^{\ast}$ that is an analytical function on $\mathbb{C}\setminus\mathbb{R}^{+}$. Let $(z_{k})_{k\in\mathbb{N}}$ be a sequence with an accumulation point in $\mathcal{D}$. Then, for each $k$, $f_{p}(z_{k})\rightarrow 0$ with probability one as $p$ tends to +$\infty$+ ∞. This implies that $f^{\ast}(z_{k})=0$ for each $k$. We finally obtain that $f^{\ast}$ is identically zero on $\mathbb{C}\setminus\mathbb{R}^{+}$. Therefore $f_{p}$ converges uniformly to zero which proves that $\frac{1}{p}\operatorname{Tr}[Q_{p}(z)U_{p}]\sim\frac{1}{p}\operatorname{Tr}[T_{p}(z)U_{p}]$ for $z\in\mathbb{C}\setminus\mathbb{R}^{+}$.
@@
[…]
@@
[…]
Since we have proved that $\frac{1}{p}\operatorname{Tr}[Q_{p}(z)U_{p}]\sim\frac{1}{p}\operatorname{Tr}[T_{p}(z)U_{p}]$ holds on $\mathbb{C}\setminus\mathbb{R}^{+}$, we obtain that for all $w\in\rho$ that $\lim_{n\to\infty,\;p/n\to c}\frac{1}{p}\mathrm{Tr}[(Q_{p}(w)-T_{p}(w))U_{p}]=0.$ Moreover, for $w\in\mathbb{C}\setminus\mathbb{R}^{+}$, we have from (A.2) that
$\frac{1}{p}\frac{|\mathrm{Tr}[(Q_{p}(w)-T_{p}(w))U_{p}]|}{|w-z|^{2}}\leq\frac{2K}{\mathrm{dist}(z,\mathbb{R}^{+})|w-z|^{2}}\leq\frac{2K}{\mathrm{dist}(\rho,\mathbb{R}^{+})|w-z|^{2}}.$
Since $\frac{2K}{\mathrm{dist}(\rho,\mathbb{R}^{+})|w-z|^{2}}$ is integrable on $\rho$, using the dominated convergence theorem, we obtain that $\frac{1}{p}\operatorname{Tr}[Q_{p}^{\prime}(z)U_{p}]\sim\frac{1}{p}\operatorname{Tr}[T^{\prime}(z)U_{p}]$ holds for $z\in\mathbb{C}\setminus\mathbb{R}^{+}$ which completes the proof of Theorem 3.1. ∎

### v4

@@ 4 Numerical experiments
In this section, we illustrate the results of this paper with numerical experiments. Numerical experiments have been conducted using a random matrix $Z_{n}=(z_{ij})_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq p\end{subarray}}$ with iid entries sampled from a centered and normalized Pareto distribution. More precisely, let $(\xi_{ij})_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq p\end{subarray}}$ be iid random variables following a Pareto distribution with parameters $(a,m)=(6,1)$, whose density is given by $f(x)=\frac{am^{a}}{x^{a+1}}=\frac{6}{x^{7}},\ x\geq 1.$ The entries of $Z_{n}$ are then defined as $z_{ij}=\frac{\sqrt{a}\,(\xi_{ij}-a)}{\sqrt{a-2}},$ so that $\mathbb{E}[z_{ij}]=0$ and $\mathbb{V}ar(z_{ij})=1$. The moments of the Pareto random variables $\xi_{ij}$ are given by $\mathbb{E}[\xi_{ij}^{k}]=\begin{cases}\dfrac{am^{k}}{a-k}=\dfrac{6}{6-k},&\text{if }k<a,\\ +\infty,&\text{otherwise}.\end{cases}$ In particular, the first five moments of $\xi_{ij}$ are finite. Consequently, the entries of $Z_{n}$ satisfy  Assumption 1.2. The test data are generated according to the same procedure. We note that we also conducted the same experiments using Gaussian random variables and obtained quantitatively identical numerical results. This observation provides additional empirical evidence in support of the universality of our theoretical findings with respect to the distribution of the design entries.
In addition, we consider several variance profiles normalized such that $\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{p}\frac{\gamma^{2}_{ij}}{p}=1.$ Apart from the constant, quasi doubly stochastic, and piecewise constant variance profiles mentioned earlier, we will use the following examples of variance profiles:
@@
A variance profile referred to as the Berlin Photo for which $\gamma_{ij}$ is equal to the value of the coordinate $(i,j)$ of the pixel of the image shown in Figure 3. The original image being of size $1250\times 850$, it is rescaled according to the values of $n$ and $p$ used in the numerical experiments thanks to the function resize from the Python module PIL.Image.
@@
The values of the predictive and training risks and their deterministic equivalent are compared for various variance profiles in Figures 4 and 4 with $\lambda$ ranging from 0.1 to 10, $n=400$ and $p=600$. The curves displayed in these figures confirm that $r_{\lambda}^{test}(X_{n})$ and $r_{\lambda}^{train}(X_{n})$ are very accurate estimators of $\hat{r}_{\lambda}^{test}(X_{n})$ and $\hat{r}_{\lambda}^{train}(X_{n})$ in high-dimension since the dashed and solid lines coincide. These variance profiles also provide curves of predictive risks having similar shapes (up to a vertical translation). The curves representing the cases of constant and quasi doubly stochastic variance profile coincide, confirming the comment made in Section 1.1 on the doubly stochastic profile and its similarities with the setting of iid data. Moreover, the minimum of $r_{\lambda}^{test}(X_{n})$ is indeed reached at the optimal value $\lambda_{\ast}$ for any variance profile as shown by Corollary 3.1.
@@
We illustrate the appearance of a double or triple descent phenomenon for several variance profiles in Figure 5 (a), (c) and (e). These figures represent the predictive risk and its approximation by a deterministic equivalent for various values of the ratio $p/n$ with $\lambda=0$. The solid lines depicts $r_{0}^{test}$ whereas the dashed lines represent $\hat{r}_{0}^{test}$. For every variance profile, the solid and dashed lines coincide which confirms that $r_{0}^{test}$ is a relevant deterministic equivalent of $\hat{r}_{0}^{test}$. For the constant and quasi doubly stochastic variance profiles, we observe the well known double descent phenomenon as illustrated by Figure 5 (a) since the curves are increasing for $p/n<1$ and decreasing for $p/n>1$. This is related to Corollary 3.3 which states that the expression of $\lim_{\lambda\rightarrow 0}r_{\lambda}^{test}(X_{n})$ depends upon the value of the ratio $p/n$ with respect to one. Nevertheless, for other variance profiles, the shape of the predictive risk can be very different from one variance profile to another, and it differs from the usual double descent. For some variance profiles a phenomenon of triple descent arises, notably for the piecewise constant and block cases, as shown in Figure 5 (c) and (e). We also remark the appearance of a quadruple descent in the case of the polynomial profile, as shown in Figure 5 (a).
@@
As already remarked at the end of Section 3, it is difficult to analytically explain these phenomenon for an arbitrary variance profile since we do not have an explicit formula for $T_{p}(z)$ in the general case. Nevertheless, as nicely discussed in [SKR+23], the double descent phenomenon is very much related to the distance from zero of the smallest non-negative eigenvalue of $\widehat{\Sigma}_{n}$. Having this eigenvalue close to zero causes double descent. In Figure 5 (b), we thus represent the smallest non-zero eigenvalue $\tau_{\min}(\widehat{\Sigma}_{n})$ of $\widehat{\Sigma}_{n}$ for many variance profiles as a function of $p/n$. It can be observed that, for all variance profiles considered, the curves mapping $p/n\mapsto\tau_{\min}(\widehat{\Sigma}_{n})$ attain their minimum at a value close to zero when $p=n$, corresponding to the interpolation threshold at which the classical double descent phenomenon occurs. This behavior is consistent with previous observations reported in [SKR+23]. A more detailed inspection of Figure 5(b) reveals a notable qualitative difference between variance profiles. In particular, for the piecewise constant and polynomial variance profiles, the smallest non-zero eigenvalue $\tau_{\min}(\widehat{\Sigma}_{n})$ remains significantly closer to zero over a wide range of aspect ratios, approximately for $p/n\in[0.25,3]$, compared to the constant or quasi–doubly stochastic cases. This behavior manifests itself as an extended plateau near zero in the corresponding spectral curves. The presence of such plateaus suggests prolonged near-singularity of the empirical covariance matrix across a broad range of over- and under-parameterized regimes. This spectral degeneracy may play a key role in the emergence of more complex risk landscapes, and in particular may be closely related to the appearance of the triple and quadruple descent phenomena observed in Figure 5(b). Similar trends are observed for the Alternated columns and Berlin photo variance profiles considered in our experiments. Although it is challenging to visually determine whether this eigenvalue reaches its minimum near the interpolation threshold $p=n$ in Figure 5(b), we include zoomed-in versions of this figure in Figure 6, which enable a more precise examination of its behavior in the neighborhood of $p/n=1$.
@@
Figure 6: Smallest non-zero eigenvalue of $\widehat{\Sigma}_{n}$ for several variance profiles with $\alpha=\sigma=1$, $n=300$ and $p$ varying from $30$ to $1800$. Left: Zoomed version of Figure 5 (f). Right: Zoomed version of Figure 5 (d).
Figure 7: Study of MNIST dataset whenever $\lambda=0$. (a) Comparison of real and synthetic data with $\alpha=\sigma=1$, $p=784$ and $n$ varying from $78$ to $1568$. The solid lines correspond to the predictive risk while the dashed line corresponds to the deterministic equivalent.(b) Heatmap of the variance of pixels for each class. These heatmaps serve as variance profiles in the mixture model created from the MNIST dataset and described in the beginning Section 4.
@@ 5 Conclusion
In this paper, we derive a deterministic equivalent of the DOF and the predictive risk of ridge (less) regression in a high-dimensional framework with a variance profile to handle the setting of non-iid data. We use RMT to determine a deterministic equivalent of the diagonal of the resolvent matrix $Q_{p}(z)$. Our work extends the study of the DOF and the predictive risk from [DW18] and [Bac24] to the case of general variance profile. It appears that also the result of [DW18] and [Bac24] still hold in the case of a quasi doubly stochastic profile. However, there are variance profiles that cause a behavior of the predictive risk of the minimum norm least square estimator that differs from the double descent phenomenon as classically observed in the case of a constant variance profile. Our numerical experiments confirm that our deterministic equivalent accurately estimates the predictive risk in high-dimension. Our results also allow to understand how assuming such a variance profile for the data influences the statistical properties of ridge regression when compared to the standard assumption of iid observations. Beyond the random-matrix analysis itself, the deterministic equivalents provided here can be viewed as a practical “calculus” for high-dimensional prediction under heterogeneous designs: they yield computable approximations for prediction error and degrees of freedom, enable principled tuning rules (oracle and GCV-type), and explain when interpolation phenomena resemble or depart from the iid benchmark depending on the variance profile. These consequences are precisely the type of statistical insights that deterministic equivalents have enabled in the iid ridge literature, and our work extends this program to non-identically distributed designs.
We hope that our approach on the use of variance profiles may lead to further research works on the statistical analysis of other estimators than ridge regression in more complex models with non-iid data.
@@
In this appendix, we give the proofs needed to establish our results. We also discuss the link between random matrices with a variance profile and the notion of $\mathcal{R}$-transform in free probability.
@@
Recall that $\frac{1}{p}\widehat{df}_{1}(\lambda)=\frac{1}{p}\operatorname{Tr}[\widehat{\Sigma}_{n}(\widehat{\Sigma}_{n}+\lambda I_{p})^{-1}]$, for $\lambda>0$. Thus by using $\hat{\mu}_{n}$ the empirical eigenvalue distribution of $\widehat{\Sigma}_{n}$ we have this new expression of the degree of freedom $\frac{1}{p}\widehat{df}_{1}(\lambda)=\int_{{\mathbb{R}}}\frac{t}{t+\lambda}d\hat{\mu}_{n}(t).$ Moreover, $\hat{\mu}_{n}$ being a probability measure one can notice from Definition 2.1 that
@@ A.2 Derivation of the deterministic equivalent the diagonal of the resolvent
This section is dedicated to prove Theorem 3.1 that aims to provide a deterministic equivalent of $\Delta[Q_{p}(z)]$ that is necessary to obtain asymptotic equivalents of the risks. The proof of this theorem is strongly inspired by results from [HLN07], and it constitutes a key element in the proof of Theorem 3.2. In order to prove Theorem 3.1, we introduce the diagonal matrices $R_{p}(z)=\underset{1\leq j\leq p}{\mathop{\rm diag}\nolimits}(R^{(j)}_{p}(z))$ defined by $R^{(j)}_{p}(z)=\frac{-1}{z\left((1+(1/n)\operatorname{Tr}[\widetilde{D}_{j}\widetilde{Q}_{n}(z)]\right)}\quad\mbox{for }z\in\mathbb{C}\setminus\mathbb{R}^{+},$ with $\widetilde{Q}_{n}(z)=\Big(\widetilde{\Sigma}_{n}-zI_{n}\Big)^{-1}$ is the resolvent matrix of $\widetilde{\Sigma}_{n}=\frac{1}{n}X_{n}X_{n}^{\top}$.
@@
Then, Lemma A.1 needs to be combined with the following one in order to prove Theorem Theorem 3.1.
@@
Hence, for each compact subset $C\subset\mathbb{C}\setminus\mathbb{R}^{+}$, $f_{p}$ is uniformly bounded on $C$, that is $|f_{p}(z)|\leq\frac{2K}{\delta_{C}}$, where $\delta_{C}$ is the distance between $C$ and $\mathbb{R}^{+}$. Then, by the normal family Theorem [Rud87][Theorem 14.6] there exists a sub-sequence $f_{p_{k}}$ which uniformly converges to $f^{\ast}$ that is an analytical function on $\mathbb{C}\setminus\mathbb{R}^{+}$. Let $(z_{k})_{k\in\mathbb{N}}$ be a sequence with an accumulation point in $\mathcal{D}$. Then, for each $k$, $f_{p}(z_{k})\rightarrow 0$ with probability one as $p$ tends to $\infty$. This implies that $f^{\ast}(z_{k})=0$ for each $k$. We finally obtain that $f^{\ast}$ is identically zero on $\mathbb{C}\setminus\mathbb{R}^{+}$. Therefore $f_{p}$ converges uniformly to zero which proves that $\frac{1}{p}\operatorname{Tr}[Q_{p}(z)U_{p}]\sim\frac{1}{p}\operatorname{Tr}[T_{p}(z)U_{p}]$ for $z\in\mathbb{C}\setminus\mathbb{R}^{+}$.
@@
$T_{p}^{\prime}(z)=\frac{1}{2\pi i}\int_{\rho}\frac{T_{p}(w)}{(w-z)^{2}}dw\quad\mathrm{and}\quad Q_{p}^{\prime}(z)=\frac{1}{2\pi i}\int_{\rho}\frac{Q_{p}(w)}{(w-z)^{2}}dw,$
@@
$\frac{1}{p}\operatorname{Tr}[Q_{p}^{\prime}(z)U_{p}]-\frac{1}{p}\operatorname{Tr}[T_{p}^{\prime}(z)U_{p}]$ $=$ $\frac{1}{p}\mathrm{Tr}\Bigg[\int_{\rho}\frac{Q_{p}(w)-T_{p}(w)}{(w-z)^{2}}dw\ U_{p}\Bigg]=\int_{\rho}\frac{1}{p}\frac{\mathrm{Tr}[(Q_{p}(w)-T_{p}(w))U_{p}]}{(w-z)^{2}}dw.$
Since we have proved that $\frac{1}{p}\operatorname{Tr}[Q_{p}(z)U_{p}]\sim\frac{1}{p}\operatorname{Tr}[T_{p}(z)U_{p}]$ holds on $\mathbb{C}\setminus\mathbb{R}^{+}$, we obtain that for all $w\in\rho$ that $\lim_{n\to\infty,\;p/n\to c}\frac{1}{p}\mathrm{Tr}[(Q_{p}(w)-T_{p}(w))U_{p}]=0.$ Moreover, for $w\in\mathbb{C}\setminus\mathbb{R}^{+}$, we have from Equation A.2 that $\frac{1}{p}\frac{|\mathrm{Tr}[(Q_{p}(w)-T_{p}(w))U_{p}]|}{|w-z|^{2}}\leq\frac{2K}{\mathrm{dist}(z,\mathbb{R}^{+})|w-z|^{2}}\leq\frac{2K}{\mathrm{dist}(\rho,\mathbb{R}^{+})|w-z|^{2}}.$ Since $\frac{2K}{\mathrm{dist}(\rho,\mathbb{R}^{+})|w-z|^{2}}$ is integrable on $\rho$, using the dominated convergence theorem, we obtain that $\frac{1}{p}\operatorname{Tr}[Q_{p}^{\prime}(z)U_{p}]\sim\frac{1}{p}\operatorname{Tr}[T^{\prime}(z)U_{p}]$ holds for $z\in\mathbb{C}\setminus\mathbb{R}^{+}$ which completes the proof of Theorem 3.1. ∎